# On the structure of length sets with maximal elasticity

### v1

@@ 2.4. Length Sets with Maximal Elasticity
Let $A\in\mathcal{B}(G)$ and suppose that $\rho(\mathsf{L}(A))=\mathsf{D}(G)/2$. Then by Lemma 3.3 there exist atoms $U_{1},.\hskip 2.84544pt.\hskip 2.84544pt.\hskip 2.84544pt,U_{m},\hskip 2.84544ptV_{1},.\hskip 2.84544pt.\hskip 2.84544pt.\hskip 2.84544pt,V_{l}\in\mathcal{A}(G)$ with $A=U_{1}\cdot\ldots\cdot U_{m}=V_{1}\cdot\ldots\cdot V_{l}$, where $|U_{i}|=\mathsf{D}(G)$ and $|V_{j}|=2$ for all $i\in[1,m]$ and $j\in[1,l]$. Lemma 5.1 ensures that $m$ is even, say $m=2k$ for some $k\in\mathbb{N}$. Then $A=U_{1}(-U_{1})\cdot\ldots\cdot U_{k}(-U_{k})$, and moreover, by applying [6, Lemma 1.4.9], we have $U_{i}=g_{i}^{n}$, where $g_{i}\in G_{0}$ for all $i\in[1,k]$. Thus, the sequence $A\in\mathcal{B}(G)$ with maximal elasticity is equal to $\prod_{i=1}^{k}g_{i}^{n}(-g_{i})^{n}$. ∎
@@
1. Let $2n\in\{4,6,10\}$. Then [8, Theorem 3.5] states that $1\notin\Delta_{\rho}(G)$, which particularly implies that $\min\mathsf{L}(A)+1\notin\mathsf{L}(A)$ for all sequences $A\in\mathcal{B}(G)$ with maximal elasticity. Suppose, for the sake of contradiction, that there exists a sequence $A\in\mathcal{B}(G)$ with maximal elasticity such that $\min\mathsf{L}(A)+1\in\mathsf{L}(A)$. By applying the same argument as in the proof of Theorem 1.2, we deduce that there exists $M\in\mathbb{N}$ such that $\mathsf{L}(A^{*}_{k})\cap[\min\mathsf{L}(A^{*}_{k}),\max\mathsf{L}(A^{*}_{k})-M]$ is an interval for all sufficiently large $k$, where $A^{*}_{k}=A^{k}$. This is equivalent to asserting that $1\in\Delta_{\rho}(G)$, which leads to a contradiction.

### v2

@@ 2.4. Length Sets with Maximal Elasticity
Let $A\in\mathcal{B}(G)$ and suppose that $\rho(\mathsf{L}(A))=\mathsf{D}(G)/2$. Then, by Lemma 3.3, there exist atoms $U_{1},.\hskip 2.84544pt.\hskip 2.84544pt.\hskip 2.84544pt,U_{m},\hskip 2.84544ptV_{1},.\hskip 2.84544pt.\hskip 2.84544pt.\hskip 2.84544pt,V_{l}\in\mathcal{A}(G)$ such that $A=U_{1}\cdot\ldots\cdot U_{m}=V_{1}\cdot\ldots\cdot V_{l}$, where $|U_{i}|=\mathsf{D}(G)$ and $|V_{j}|=2$ for all $i\in[1,m]$ and $j\in[1,l]$. Lemma 5.1 ensures that $m$ is even, say $m=2k$ for some $k\in\mathbb{N}$. By applying [6, Lemma 1.4.9], we obtain that $U_{i}=g_{i}^{n}$, where $g_{i}\in G_{0}$ for all $i\in[1,k]$. Thus, we conclude that the sequence $A\in\mathcal{B}(G)$ with maximal elasticity is equal to $\prod_{i=1}^{k}U_{i}(-U_{i})=\prod_{i=1}^{k}g_{i}^{n}(-g_{i})^{n}$ (after reindexing if necessary). ∎
@@
1. Let $2n\in\{4,6,10\}$. Then [8, Theorem 3.5] states that $1\notin\Delta_{\rho}(G)$. We claim that this implies that $\min\mathsf{L}(A)+1\notin\mathsf{L}(A)$ for every sequences $A\in\mathcal{B}(G)$ with maximal elasticity. Suppose, seeking a contradiction, that there exists a sequence $A\in\mathcal{B}(G)$ with maximal elasticity such that $\min\mathsf{L}(A)+1\in\mathsf{L}(A)$. By applying the same argument as in the proof of Theorem 1.2, we deduce that there exists $M\in\mathbb{N}$ such that $\mathsf{L}(A^{*}_{k})\cap[\min\mathsf{L}(A^{*}_{k}),\max\mathsf{L}(A^{*}_{k})-M]$ is an interval for all sufficiently large $k$, where $A^{*}_{k}=A^{k}$. This is equivalent to asserting that $1\in\Delta_{\rho}(G)$, which leads to a contradiction.
@@
Conflict of interest statement. The author states that there is no conflict of interest.